# AVA: Attentive VLM Agent for Mastering StarCraft II

### v3

@@ Appendix A Map Details
A.1 Single Player Maps (No Abilities)
@@
6reaper_vs8zealot_vlm_attention Player: 6 Reapers
@@
8marine_1medvac_vs_2tank Player: 8 Marines, 1 Medivac
@@
8marine_2tank_vs_zerglings _banelings_vlm_attention Player: 8 Marines, 2 Siege Tanks
@@
2bc1prism_vs_8m_vlm_attention Player: 8 Marines
@@
2s3z_vlm_attention Player: 2 Stalkers, 3 Zealots
@@
vlm_attention_1 Player: 1 Zealot, 1 Immortal, 1 Archon, 1 Stalker, 1 Phoenix
@@
A.2 Single Player Maps (With Abilities)
@@
ability_map_8marine _3marauder_1medivac_1tank Player: 8 Marines, 3 Marauders, 1 Medivac, 1 Siege Tank
@@
ability_8stalker_vs _8marine_3marauder _1medivac_tank_map Player: 8 Stalkers
@@
ability_map_8marine _3marauder _1medivac_1tank _vs_5stalker_2colossis Player: 8 Marines, 3 Marauders, 1 Medivac, 1 Siege Tank
@@
pvz_task6_level3 Player: 12 Stalkers, 1 Archon, 4 Sentries, 6 High Templars
@@
A.3 Two Player Maps (No Abilities)
@@
MMM_vlm_attention_two_players Player 1: 8 Marines, 3 Marauders, 1 Medivac
@@
vlm_attention_1_two_players Player 1: 1 Zealot, 1 Immortal, 1 Archon, 1 Stalker, 1 Phoenix
@@
vlm_attention_2_terran_vs _terran_two_players Player 1: 1 Marine, 1 Marauder, 1 Reaper, 1 Hellbat, 1 Medivac, 1 Viking (Assault), 1 Ghost, 1 Banshee
@@
A.4 Two Player Maps (With Abilities)
@@
ability_7stalker_vs _11marine_1medivac _1tank_map_2_players Player 1: 7 Stalkers
@@
ability_8stalker_vs _8marine_3marauder _1medivac_tank_map_2_players Player 1: 8 Stalkers
@@
ability_map_8marine _3marauder_1medivac _1tank_2_players Player 1: 8 Marines, 3 Marauders, 1 Medivac, 1 Siege Tank
@@
(a) Original RGB observation of battlefield situation
(b) Annotated unit positions with unit IDs and health status
(c) AVA’s strategic analysis highlighting prioritized targets and optimal attack vectors
Figure 4: Visualization of AVA’s decision-making process in a complex micro-management scenario. The system processes the raw visual input (a), analyzes unit positions and states (b), and makes tactical decisions (c). In this case, AVA identified Zergling_52 (Tag: 54) as a priority target due to its strategic position at [2,1], where attacking it would maximize area-of-effect damage to nearby clustered units. This decision demonstrates the system’s ability to not only identify low-health targets (5/35 HP) but also recognize opportunities for efficient damage distribution through Colossi’s line damage mechanic. Supporting this decision, the system also identified Zergling_1 (Tag: 3) and Zergling_2 (Tag: 4) as secondary priority targets due to their threatening positions at [1,1] and [0,1] ,respectively, enabling a comprehensive control strategy that combines focus fire with positional advantage.
@@
(d) Marine formation adjustment
(e) Coordinated focus fire execution
(f) Optimized Marine positioning
(g) Secondary target engagement
(h) Maintained spread formation
(i) Final engagement phase
Figure 5: Demonstration of AVA’s micro-management decision process in a Marine/Tank vs Baneling/Zergling scenario. The system implements a three-stage strategy: (1) Initial threat assessment prioritizing Banelings, (2) Coordinated unit positioning with Marines at safe distances ([0,1] and [0,2]) while maintaining focus fire capability, and (3) Sequential target elimination following the identified priority order of Banelings first, then supporting Zerglings. This showcases the system’s ability to combine Focus Fire, Formation Control, and Shield/Health Management skills in a complex tactical situation.
(a) Initial state analysis with two Siege Tanks identified as primary threats
(b) Execution of three-way split formation (3-3-2) with Medivac support
Figure 6: Visualization of AVA’s Marine split micro-management against Siege Tanks. The system first analyzes the battlefield state (a), identifying two Siege Tanks (Tag: 10, 11, Health: 160/160) as critical threats. Then it executes a sophisticated tactical response (b) by implementing a three-way Marine split formation (3-3-2) to minimize splash damage vulnerability while maintaining effective focus fire. The coordinated execution combines multiple micro-skills: Focus Fire for optimal damage distribution, Formation Control through strategic three-way positioning, and Health Management with Medivac support positioned at [0,2]. This demonstrates the system’s ability to balance multiple tactical objectives while maintaining combat effectiveness against superior firepower.
@@
Figure 7: AVA’s multi-type unit coordination in a complex Protoss vs Terran engagement. The system first performs threat assessment (a) by identifying critical targe,ts including a low-health Viking Assault (11/125 HP), energy-rich Ghost (56 energy), and support units like Medivac. Based on this analysis, it executes a coordinated attack plan (b) where each unit is assigned optimal targets: Zealot engages the weakened Viking, Phoenix provides air superiority against Medivac, Immortal focuses on armored targets, while the Archon maintains a strategic position at [7,7] for battlefield control. The decision-making process demonstrates VLM’s understanding of unit-specific attributes (health states, energy levels, armor types) and tactical synergies in a mixed-unit combat scenario. This showcases the system’s ability to handle complex micro-management tasks involving heterogeneous unit compositions without requiring explicit training for each unit combination.
Figure 8: Colossis stand in corner

### v5

@@ Appendix D Map Details
D.1 Single Player Maps (No Abilities)
@@
6r_vs_8z Player: 6 Reapers
@@
8m1mv_vs_2st Player: 8 Marines, 1 Medivac
@@
8m2st_vs_35zg4b Player: 8 Marines, 2 Siege Tanks
@@
8m_vs_2pc1wp Player: 8 Marines
@@
2s3z Player: 2 Stalkers, 3 Zealots
@@
mixed_units Player: 1 Zealot, 1 Immortal, 1 Archon, 1 Stalker, 1 Phoenix
@@
D.2 Single Player Maps (With Abilities)
@@
8m3mr1mv1st_mirror Player: 8 Marines, 3 Marauders, 1 Medivac, 1 Siege Tank
@@
8s_vs_8m3mr1mv1st Player: 8 Stalkers
@@
8m3mr1mv1st_vs_5s2c Player: 8 Marines, 3 Marauders, 1 Medivac, 1 Siege Tank
@@
pvz_ht Player: 12 Stalkers, 1 Archon, 4 Sentries, 6 High Templars
@@
D.3 Two Player Maps (No Abilities)
@@
MMM_vs_MMM Player 1: 8 Marines, 3 Marauders, 1 Medivac
@@
mixed_units_pvp Player 1: 1 Zealot, 1 Immortal, 1 Archon, 1 Stalker, 1 Phoenix
@@
terran_mirror Player 1: 1 Marine, 1 Marauder, 1 Reaper, 1 Hellbat, 1 Medivac, 1 Viking (Assault), 1 Ghost, 1 Banshee
@@
D.4 Two Player Maps (With Abilities)
@@
7s_vs_11m1mv1st Player 1: 7 Stalkers
@@
8s_vs_8m3mr1mv1st_pvp Player 1: 8 Stalkers
@@
8m3mr1mv1st_mirror_pvp Player 1: 8 Marines, 3 Marauders, 1 Medivac, 1 Siege Tank
@@
Appendix E Evaluation Metrics
We define the three metrics used in the human evaluation of AVA and MARL agents, each rated on a 1–5 scale:
  • Game Bug Exploitation: Measures whether the agent exploits game bugs, particularly vulnerabilities in SMAC’s built-in AI, which uses a flawed strategy of attacking only the enemy’s spawn point and stopping if the enemy moves out of range or beyond attack distance (1 = frequent exploitation, 5 = no exploitation).
  • Reasoning Coherence: Evaluates whether the agent’s decisions are logical, incorporating StarCraft II game knowledge (e.g., unit matchups) and operational skills (e.g., positioning, targeting) (1 = illogical, 5 = perfect logic).
  • Human Similarity: Assesses how closely the agent’s strategies resemble human play, including techniques like hit-and-run tactics and multi-unit coordination (e.g., combined-arms strategies) (1 = unlike human, 5 = completely human-like).
Appendix F Case of Study
Figure 4: Original RGB observation of battlefield situation in the Colossi vs Zerglings scenario.
Figure 5: Annotated unit positions with unit IDs and health status.
Figures 4 and 5 illustrate the initial stages of AVA’s decision-making process. The system begins by processing the raw RGB battlefield observation, then identifies and annotates individual units with their respective IDs and health status. This visual processing stage forms the foundation for subsequent tactical analysis.
Figure 6: AVA’s strategic analysis highlighting prioritized targets and optimal attack vectors.
Figure 6 demonstrates AVA’s strategic decision-making capabilities. In this complex micro-management scenario, AVA identified Zergling_52 (Tag: 54) as a priority target due to its strategic position at [2,1], where attacking it would maximize area-of-effect damage to nearby clustered units. This decision demonstrates the system’s ability to not only identify low-health targets (5/35 HP) but also recognize opportunities for efficient damage distribution through Colossi’s line damage mechanic. Supporting this decision, the system also identified Zergling_1 (Tag: 3) and Zergling_2 (Tag: 4) as secondary priority targets due to their threatening positions at [1,1] and [0,1] respectively, enabling a comprehensive control strategy that combines focus fire with positional advantage.
@@ Appendix F Case of Study
Figure 7: Stage 1: AVA’s battlefield analysis and threat assessment in Marine/Tank vs Baneling/Zergling engagement.
(a) Marine formation adjustment
(b) Coordinated focus fire execution
(c) Optimized Marine positioning
Figure 8: Stage 2: Tactical positioning and focus fire coordination on priority targets.
(a) Secondary target engagement
(b) Maintained spread formation
(c) Final engagement phase
Figure 9: Stage 3: Sequential target elimination while maintaining strategic formation.
The tactical execution depicted in Figures 7, 8, and 9 showcases AVA’s sophisticated decision-making processes that emerge without explicit training. The system first performs battlefield analysis, identifying Banelings as primary threats due to their splash damage potential against clustered units. It then implements a coordinated response by strategically positioning Marines at safe distances while maintaining focus fire capabilities. Throughout the engagement, AVA demonstrates multiple micro-skills simultaneously: prioritized target selection, formation control, and adaptive positioning. This behavior closely resembles human expert gameplay strategies, highlighting AVA’s ability to leverage VLM reasoning for complex tactical decision-making that would typically require extensive reinforcement or imitation learning in traditional approaches.
@@
Figure 10: Multi-type unit coordination in Protoss vs Terran engagement, showing AVA’s strategic targeting based on unit attributes and tactical synergies.
Figure 10 illustrates AVA’s ability to coordinate heterogeneous unit compositions. In the initial analysis phase (a), the system identifies critical targets including a low-health Viking Assault (11/125 HP), an energy-rich Ghost (56 energy), and support units like Medivac. Based on this assessment, it executes a coordinated attack plan (b) where each unit is assigned optimal targets: Zealot engages the weakened Viking, Phoenix provides air superiority against Medivac, Immortal focuses on armored targets, while the Archon maintains a strategic position for battlefield control. This demonstrates VLM’s understanding of unit-specific attributes (health states, energy levels, armor types) and tactical synergies in mixed-unit scenarios without requiring explicit training.
AVA demonstrates robust performance in scenarios requiring strategic target selection and basic coordination but encounters challenges with complex micro-management tasks requiring precise ability timing (as in 2s_vs_1sc_vlm_priority) or sophisticated terrain exploitation (as in 2c_vs_64zg_vlm_priority, Figure 11). Through systematic analysis, we identified three primary limitations: (1) inconsistent spatial understanding in dense unit formations, (2) challenges in maintaining temporal consistency during high-frequency decision cycles, and (3) occasional misapplication of domain knowledge in rapidly evolving scenarios with high unit count.
Figure 11: Tactical terrain exploitation: Colossi positioned in corner location to maximize attack range while minimizing exposure to enemy units.